a plane lands at a speed of 68 m/s and slows down at a rate 0f 4 m/s^2. How much runway is needed to stop the plane?


1
Expert's answer
2021-02-15T17:03:34-0500

Answer

By applying third equation of motion

"v^2=u^2+2as"

When plane will stop then

"v=0m\/s"

Acceleration

"a=-4m\/s^2"

Initial velocity

"u=68m\/s"


So by putting value

"0^2=68^2-2(4) (s) \\s=\frac{68^2}{8}\\s=578m"

So 578m runway is required.